According to a recent​ study, the carapace length for adult males of a certain species of tarantula are normally distributed with a mean of 18.66 mm and a standard deviation of 1.55 mm. Find the percentage of the tarantulas that have a carapace length exceeding 20 mm.
Answer #1

Define random variable X: Carapace length for adult males of a certain species of tarantula

X is normally distributed with mean = \mu = 18.66 and standard deviation = \sigma = 1.55

Here we have to find P(X > 20)

P(X>20)=P(\frac{X-\mu}{\sigma}>\frac{20-\mu}{\sigma})

=P(z>\frac{20-18.66}{1.55}) where z is standard normal variable

=P(z>\frac{1.34}{1.55})

= P(z > 0.86) (Round to 2 decimal)

= 1 - P(z < 0.86)

= 1 - 0.8051 (From statistical table of z values)

= 0.1949

19.49 % of the tarantulas that have a carapace length exceeding 20 mm.